Add. Simplify completely. Write as a mixed number if possible.
15 4/15 + 5 5/6

Answer by stanbon(75887)

Write as a mixed number if possible.
15 4/15 + 5 5/6
----
(225+4)/15 + (30+5)/6
----
lcd = 30
---
= (2*229)/30 + (5*35)/30
---
= (458+175)/30
---
= (633/30)
---
= 211/10
= 21 1/10
==============
Cheers,
Stan H.